OpenStudy (anonymous):

The sides of a triangle are x, x+1, 2x-1 and its area is x root 10 .What is the value of x?

OpenStudy (anonymous):

So if the sides = x>-2 Then the area would be that ^10 right?

OpenStudy (anonymous):

cn u xplain??

OpenStudy (raden):

Heron's formula is work area = sqrt(s(s-a)(s-b)(s-c)) with s = (perimeter of triangle)/2 a,b,c are the sides triangle

OpenStudy (raden):

first, determine the perimeter of that triangle perimeter = a + b + c = x + x+1 + 2x - 1 = 4x so, s = perimeter/2 = 4x/2 = 2x

OpenStudy (raden):

now, use the Heron's formula above area = sqrt(s(s-a)(s-b)(s-c)) known area of that triangle is xsqrt(10), so we have the equation sqrt(s(s-a)(s-b)(s-c)) = x sqrt(10)

OpenStudy (raden):

distribute the values of s = 2x, a = x, b = x+1, and c = 2x -1 into equation above : sqrt(s(s-a)(s-b)(s-c)) = x sqrt(10) sqrt(2x(2x- x)(2x - (x+1))(2x - (2x-1)) = x sqrt(10) simplify, we get sqrt(2x * x * (x-1) * 1) = x sqrt(10) sqrt(x^2 * 2(x-1)) = x sqrt(10) sqrt(x^2) * sqrt(2(x-1)) = x sqrt(10) x sqrt(2(x-1)) = x sqrt(10)

OpenStudy (raden):

just take identicaly, the value in radical must be same. so 2(x-1) = 10 divided by 2 on both sides, we get x-1 = 10/2 x-1 = 5 add by 1 to both sides, get x = 5 + 1 = 6

OpenStudy (raden):

does that make sense, @juhaINA ?

OpenStudy (anonymous):

yah..!!:)

OpenStudy (raden):

ok, glad can help you :)

OpenStudy (anonymous):

i hv a question.

OpenStudy (raden):

what is that ?

OpenStudy (anonymous):

in d formula:s-a,s-b,s-c.v don hv to do dat in this question?

OpenStudy (raden):

well it is already be clear for you, right. this site might useful :) http://www.mathsisfun.com/geometry/herons-formula.html

OpenStudy (anonymous):

k..thnx

OpenStudy (raden):

you're welcome